# Just Trial Once: Ongoing Causal Validation of Machine Learning Models

### v1

@@ 1 Introduction
Figure 1: (1(a)) The goal of this paper is to predict the causal impact of deploying a new model $\pi_{e}$, given data from a cluster randomized trial that randomizes sets of users (e.g., hospitals) to one of a set of trial models that does not include $\pi_{e}$. (1(b)) The first challenge: Relevant outcomes (e.g., of patients) are not only influenced by model outputs, but also by how users actually respond to the outputs of model-based decision-support, which may itself be affected by the perceived reliability / performance of the model. (1(c)) The second challenge: There may exist some subset of cases (e.g., patients) for whom we never observe certain model outputs, making it impossible to give precise predictions for outcomes of patients in that group, denoted by the question mark.
@@
  • We provide assumptions (Assumptions 2.1, 3.1, 3.2 and 3.3) under which we derive bounds (Assumption 3.3) on the effect of deploying a new ML model, given data from a prior RCT. Our bounds are tight, and cannot be improved without further assumptions (Assumption 3.3).
  • We provide a simple estimator for these bounds and a procedure for generating asymptotically valid confidence intervals (Assumption 3.3). We also show that our core assumptions can be falsified via hypothesis tests constructed from RCT data trialing multiple models (Assumptions 3.1 and 3.2).
  • We provide recommendations for pre-trial design and post-trial analysis in light of our results (Section 4), and demonstrate in a simulation study (Section 5) that our bounds provide a more informative tool to select among model updates, as compared to using the raw performance (e.g., accuracy) of updated models.
Related literature: Our work is related to off-policy policy evaluation in causal inference and reinforcement learning [Uehara et al., 2022]. An ML-DST can be viewed as a deterministic policy that chooses actions (i.e., predictions or alerts to raise) based on context (i.e., inputs to the model) with the goal of obtaining some reward (i.e., positively influencing outcomes for patients). Two critical distinctions arise in our work versus the standard setting: First, the policies that are present in retrospective data (in our case, from a trial) are deterministic rather than random, leading to violation of the common assumption that, for a given context, there is a positive probability of seeing any action. Second, we allow for the fact that actions taken for one patient can influence outcomes for other patients. Our work is also related to causal evaluations of AI-assisted decision-making in criminal justice settings [Imai et al., 2023, Ben-Michael et al., 2024], but our goal differs: Rather than evaluating the impact of AI-assistance on the accuracy of (observable) predictions made by a human, we are interested in the total effect on downstream outcomes.
@@ 2 Model and Problem Setup
Notation: Note that in the rest of this paper, we use the terms model and policy interchangeably. We use upper case letters $X$ to denote a random variable, calligraphic font to denote the space of possible values $\mathcal{X}$, and lower-case to denote a specific realization of a random variable $x$. We assume that the causal structure of an RCT is modeled by a directed acyclic graph (DAG) $\mathcal{G}$ over a set of vertices ${\bf V}=\{A,Y,D,X,\Pi,M\}$, where $A\in\mathcal{A}$ represents the output (or “action”) of the deployed model, $Y\in\mathbb{R}$ represents an outcome of interest, $D\in\mathcal{D}$ represents the cluster to which a user is assigned, $X\in\mathcal{X}$ represents covariates used as inputs to the ML model, $\Pi\in\varPi$ represents the specific ML model that was deployed, and $M\in\mathbb{R}$ represents model performance, which we represent as a real number. We assume that model performance is computable for any model via some functional $f_{M}(\pi)$ (e.g., the accuracy, precision, recall, sensitivity, specificity, or some combination, computed on a held-out dataset where $\pi(X)$ is considered the model prediction). We also use the indicator function $\mathbf{1}\{S\}$ which is equal to $1$ if the event $S$ is true, and 00 otherwise.
@@
Figure 2: The directed acyclic graph (DAG) $\mathcal{G}$ depicting the causal relationships in our problem setup (Assumption 2.1). We draw circles around nodes $\Pi,A,M$ to represent that these variables are deterministic given their parents.
@@
For concreteness in the remainder, we will primarily use the language of healthcare applications (e.g., patients, clinical outcomes, etc). Our assumed causal structure can be represented by the structural causal model (SCM) that we define below, which is consistent with the DAG shown in Fig. 2.
@@ 3 Identification and Bounds
For simplicity, suppose that the model alerts based on thresholding a pre-defined risk score $r(x)$ that is a function of vital signs (e.g., systolic blood pressure, respiratory rate, etc). Suppose that an initial RCT assigns patients to a control arm, $D=0$ where $\pi_{0}=0$ (alerts are never raised), and a treatment arm, $D=1$ where the model raises alerts using the threshold $t$, i.e., $\pi_{1}(x)\coloneqq\mathbf{1}\{r(x)>t\}$. Suppose we want to use this RCT data to evaluate the impact of two alternative models with a lower and higher threshold respectively, $\pi_{l}(x)\coloneqq\mathbf{1}\{r(x)>t_{l}\}$ and $\pi_{u}(x)\coloneqq\mathbf{1}\{r(x)>t_{u}\}$ where $t_{u}>t>t_{l}$. Fig. 3 visually demonstrates the challenges of this task for evaluating $\pi_{l}$, since we never observe alerts for patients with $r(x)\in[t_{l},t]$.
Figure 3: Illustration from Example 1 below, demonstrating the challenge of our task. We can use data from both the control arm $\pi_{0}$ (green) and $\pi_{1}$ (blue) to infer patient outcomes when $\pi_{l}$ (orange) does not raise alerts for patients with risk scores less than $t_{l}$. Next, we can use data from $\pi_{1}$ (blue) to infer patient outcomes when $\pi_{l}$ raises alerts for patients with risk scores greater than $t$. However, we do not know what patient outcomes are when $\pi_{l}$ raises alerts for patients with risk scores between $t_{l}$ and $t$.
@@
$Y(A=a,M=m_{i})\leq Y(A=a,M=m_{j}).$
Assumption 3.1 says that improvements in model performance do not harm patient outcomes, given a fixed action. For instance, in the context of Example 1, we might expect that for a given patient, having an alarm raised by a high-performance model would not lead to worse outcome than if that alert had been raised by a model with frequent false alarms. Note that this assumption is stated with a fixed action $A=a$, and does not imply that improving performance alone is guaranteed to improve outcomes: A change in model performance is generally associated with a change in outputs, which may have its own effect on outcomes. In Section 5, we give a case where improved overall performance (accuracy) is associated with worse outcomes.
@@
implies that Assumption 3.1 is false. The proof for Assumption 3.1, along with all other proofs, is given in Appendix B. While it is not possible to guarantee that Assumption 3.1 is true in general (over all possible models), it has observable implications in an RCT that we can check. In particular, this result suggests a simple hypothesis test that we can use to falsify Assumption 3.1: compare two empirical means in the data and check if outcomes are lower under $\pi_{2}$ than under $\pi_{1}$ on those cases where $\pi_{1}$ and $\pi_{2}$ agree on their actions.
In our discussions thus far, we have considered the control arm to be just another policy. However, this framework creates practical difficulties when considering the model performance of a control arm. For instance, suppose the relevant metric for model performance is the false positive rate (as in Example 1); then, $M$ is not clearly defined because the control arm never raises alerts. Alternatively, if the relevant metric were model accuracy under no deferral (as in Example 2), then the performance of the control arm would be similarly undefined. One way of resolving this tension is to presume the existence of a “neutral” action (e.g., not raising an alert, or deferring to clinicians).
@@
Assumption 3.2 is a sufficient condition for leveraging data from the control arm of an RCT in our setting, and implies that when a model output is “neutral” (e.g., no alert in Example 1, or deferral in Example 2), decision-makers act as they would if no model were deployed [^3] Note that, while all of our results make use of Assumption 3.2, they can also be re-written to hold if Assumption 3.2 is false, by re-defining $a_{0}$ as some placeholder model output that is never observed under any model (including $\pi_{e}$), such that indicators like $\mathbf{1}\{\pi_{e}(x)=a_{0}\}$ are always zero..
@@
Consider the evaluation policy $\pi_{u}(x)=\mathbf{1}\{r(x)>t_{u}\}$ where the threshold for alerting $t_{u}>t$ is greater than the threshold used in the original trial and where the performance of $\pi_{u}(x)$ (e.g., the precision) is greater than that of the original trialed policy $\pi_{1}$. In this scenario, under Assumptions 3.1 and 3.2, we can intuitively infer a lower bound on the policy value of $\pi_{u}$ using the outcomes of both (a) patients with $r(x)\leq t_{u}$ who did not receive alerts in the trial (either because they were in the control arm or because $\pi_{1}$ did not raise alerts), and (b) those patients with $r(x)>t_{u}$ who did receive alerts under $\pi_{1}$.
While these assumptions are sufficient in some scenarios, they do not yield meaningful bounds when a new policy takes actions (i.e., a new model produces outputs) on a given case that were never seen for similar cases in the RCT.
@@
Consider the evaluation policy $\pi_{l}(x)=\mathbf{1}\{r(x)>t_{l}\}$ where the threshold for alerting $t_{l}<t$ is less than the threshold used in the original trial. Regardless of the performance of $\pi_{l}$ in this scenario, even under Assumptions 3.1 and 3.2 we have no way to infer outcomes under $\pi_{l}$ for the individuals where $r(x)\in[t_{l},t]$. These correspond to a set of “never alerted” individuals where $\pi_{l}$ raises an alert, but where neither the control arm nor the trialed policy $\pi_{1}$ raised an alert.
@@
We will shortly present our main result: Under our data-generating process (Assumption 2.1) and the assumptions above (Assumptions 3.1, 3.2 and 3.3), we can provide tight bounds on expected outcomes under any proposed model $\pi_{e}$. First, however, we will define some useful notation for conveying our results, which builds upon the intuition above.
@@
[Partitions]definitionPartitionsDef For each value of $x$, we first define the set of trialed models (possibly none) that agree with $\pi_{e}(x)$ [^4] All these partitions are defined with respect to the model $\pi_{e}$, and could be written more precisely with $\pi_{e}$ as an argument (e.g., $\mathcal{S}_{*}(x;\pi_{e})$, $\mathcal{X}_{=}(\pi_{e})$) but we suppress this notation for simplicity. and subsets of this set based on the performance characteristics of those trialed models.
$\mathcal{S}_{*}(x)$$\coloneqq\{i\in D\mid\pi_{i}(x)=\pi_{e}(x)\}$
$\mathcal{S}_{=}(x)$$\coloneqq\{i\in\mathcal{S}_{*}\mid f_{M}(\pi_{i})=f_{M}(\pi_{e})\}$
$\mathcal{S}_{<}(x)$$\coloneqq\{i\in\mathcal{S}_{*}\mid f_{M}(\pi_{i})<f_{M}(\pi_{e})\}$
$\mathcal{S}_{>}(x)$$\coloneqq\{i\in\mathcal{S}_{*}\mid f_{M}(\pi_{i})>f_{M}(\pi_{e})\}$
and we further define subsets of $\mathcal{S}_{<},\mathcal{S}_{>}$ where we only index the next-best or next-worst performing model [^5] We only use these sets when $\mathcal{S}_{<},\mathcal{S}_{>}$ are non-empty, but we additionally use the convention that $\operatorname*{arg\,min}_{i\in\varnothing}(f(i))=\varnothing$.
$\tilde{\mathcal{S}}_{<}(x)$ $\coloneqq\operatorname*{arg\,max}_{i\in\mathcal{S}_{<}(x)}f_{M}(\pi_{i}),$ $\tilde{\mathcal{S}}_{>}(x)$ $\coloneqq\operatorname*{arg\,min}_{i\in\mathcal{S}_{>}(x)}f_{M}(\pi_{i}).$
Using these sets, we define several subsets of $\mathcal{X}$, where there exist trialed models that do or do not agree with $\pi_{e}$ and have equal/better/worse performance.
$\mathcal{X}_{*}$$\coloneqq\{x\in\mathcal{X}\mid S_{*}(x)\neq\varnothing\}$
$\mathcal{X}_{=}$$\coloneqq\{x\in\mathcal{X}\mid\mathcal{S}_{=}(x)\neq\varnothing\}$
$\mathcal{X}_{<}$$\coloneqq\{x\in\mathcal{X}\mid\mathcal{S}_{<}(x)\neq\varnothing\}\setminus\mathcal{X}_{=},$ $\mathcal{X}_{\not\leq}$ $=\mathcal{X}\setminus\{\mathcal{X}_{=},\mathcal{X}_{<}\}$
$\mathcal{X}_{>}$$\coloneqq\{x\in\mathcal{X}\mid\mathcal{S}_{>}(x)\neq\varnothing\}\setminus\mathcal{X}_{=},$ $\mathcal{X}_{\not\geq}$ $=\mathcal{X}\setminus\{\mathcal{X}_{=},\mathcal{X}_{>}\}$
where $\mathcal{X}_{<},\mathcal{X}_{>}$ are defined to exclude any overlap with $\mathcal{X}_{=}$.
Using Assumption 3.3, we can give precise upper and lower bounds on the performance of a new model $\pi_{e}$. {restatable}theoremBoundsWithControl Given the data generating process in Assumption 2.1, and under Assumptions 3.3, 3.1 and 3.2, the policy value of a model / policy $\pi_{e}$ is bounded as

### v2

@@ 1 Introduction
Figure 1: (1(a)) The goal of this paper is to predict the causal impact of deploying a new model $\pi_{e}$, given data from a cluster randomized trial that randomizes sets of users (e.g., hospitals) to one of a set of trial models that does not include $\pi_{e}$. (1(b)) The first challenge: Relevant outcomes (e.g., of patients) are not only influenced by model outputs, but also by how users actually respond to the outputs of model-based decision-support, which may itself be affected by the perceived reliability / performance of the model. (1(c)) The second challenge: There may exist some subset of cases (e.g., patients) for whom we never observe certain model outputs, making it impossible to give precise predictions for outcomes of patients in that group. In this example, $\pi_{1}$ alerts for patients with high blood pressure, and $\pi_{2}$ alerts for patients with high heart rate; thus, patients with high blood pressure and high heart rate receive an alert from both $\pi_{1}$ and $\pi_{2}$. However, patients in the white region – those who have both low to medium blood pressure and heart rate – never receive an alert from either trialed model.
@@
  • We provide assumptions (Assumptions 2.1, 3.1, 3.2 and 3.3) under which we derive bounds (Remark 3.1) on the effect of deploying a new ML model, given data from a prior RCT. Our bounds are tight, and cannot be improved without further assumptions (Remark 3.1).
  • We provide a simple estimator for these bounds and a procedure for generating asymptotically valid confidence intervals (Remark 3.1). We also show that our core assumptions can be falsified via hypothesis tests constructed from RCT data trialing multiple models (Assumptions 3.1 and 3.2).
  • We provide recommendations for pre-trial design and post-trial analysis in light of our results (Section 4), and demonstrate in a simulation study (Section 5) that our bounds provide a more informative tool to select among model updates as compared to using the raw performance (e.g., accuracy) of updated models.
Related literature: Our work is related to off-policy policy evaluation in causal inference and reinforcement learning [Uehara et al., 2022]. An ML-DST can be viewed as a deterministic policy that chooses actions (i.e., predictions or alerts to raise) based on context (i.e., inputs to the model) with the goal of obtaining some reward (i.e., positively influencing outcomes for patients). Two critical distinctions arise in our work versus the standard setting: First, the policies that are present in retrospective data (in our case, from a trial) are deterministic rather than random, leading to violation of the common assumption that, for a given context, there is a positive probability of seeing any action. Second, we allow for the fact that actions taken for one patient can influence outcomes for other patients.
Our work is also related to causal evaluations of AI-assisted decision-making in criminal justice settings [Imai et al., 2023, Ben-Michael et al., 2024, 2025], but our goal differs: Rather than evaluating the impact of AI-assistance on the accuracy of (observable) predictions made by a human, we are interested in the total effect of model deployment on downstream outcomes. Finally, our work is connected to the study of causal transportability where the goal is to infer the effect of a known intervention from an RCT onto a new target population where randomization is difficult, expensive, or impossible [Pearl and Bareinboim, 2011, Stuart et al., 2011]. This is similar to our setting as we are also attempting to draw inference from RCT data. However, our problem differs in that we would like to infer the effect of an unseen intervention (i.e., a new model with no historical trial data) on the same population as in the original RCT.
@@ 2 Model and Problem Setup
Notation: In the rest of this paper, we use the terms model and policy interchangeably. We use upper case letters $X$ to denote a random variable, calligraphic font $\mathcal{X}$ to denote the space of possible values, and lower-case letters $x$ to denote a specific realization of a random variable. We assume that the causal structure of an RCT is modeled by a directed acyclic graph (DAG) $\mathcal{G}$ over a set of vertices ${\bf V}=\{A,Y,D,X,\Pi,M\}$, where $A\in\mathcal{A}$ represents the output (or “action”) of the deployed model, $Y\in\mathbb{R}$ represents an outcome of interest, $D\in\mathcal{D}$ represents the cluster to which a user is assigned, $X\in\mathcal{X}$ represents covariates used as inputs to the ML model, $\Pi\in\varPi$ represents the specific ML model that was deployed, and $M\in\mathbb{R}$ represents model performance, which we represent as a real number. We assume that model performance is computable for any model via some functional $f_{M}(\pi)$ (e.g., the accuracy, precision, recall, sensitivity, specificity, or some combination, computed on a held-out dataset where $\pi(X)$ is considered the model prediction). We also use the indicator function $\mathbf{1}\{S\}$ that is equal to $1$ if the event $S$ is true, and 00 otherwise.
@@
Figure 2: The directed acyclic graph (DAG) $\mathcal{G}$ depicting the causal relationships in our problem setup (Assumption 2.1). We draw circles around the nodes $\Pi,A,$ and $M$ to represent that these variables are deterministic given their parents (the nodes that have a direct edge to them).
@@
For concreteness in the remainder of this paper, we will primarily use the language of healthcare applications (e.g., patients, likelihood of disease onset, clinical outcomes, etc). Our assumed causal structure can be represented by the structural causal model (SCM) [Pearl, 2009] that we define below, which is consistent with the DAG shown in Fig. 2.
@@ 3 Identification and Bounds
Suppose that the trialed model alerts based on thresholding a pre-defined risk score $r(x)$ that is a function of vital signs (e.g., systolic blood pressure, respiratory rate, etc). Suppose that an initial RCT assigns patients to a control arm, $D=0$ where $\pi_{0}=0$ (alerts are never raised), and a treatment arm, $D=1$ where the model raises alerts using the threshold $T^{*}$, i.e., $\pi_{1}(x)\coloneqq\mathbf{1}\{r(x)>T^{*}\}$. Suppose we want to use this RCT data to evaluate the impact of an alternative model with a lower threshold, $\pi_{l}(x)\coloneqq\mathbf{1}\{r(x)>T^{l}\}$ where $T^{l}<T^{*}$. Fig. 3 visually demonstrates the challenges of this inference task for $\pi_{l}$ as we never observe alerts for patients with $r(x)\in[T^{l},T^{*}]$.
Figure 3: Illustration from Example 1, demonstrating the challenge of our task. We can use data from both the control arm, that does not raise alerts, and the treatment arm, that raises alerts for patients with risk scores greater than $T^{*}$, to infer patient outcomes when $\pi_{l}$ does not raise alerts for patients with risk scores less than $T^{l}$. Next, we can use data from $\pi_{1}$ to infer patient outcomes when $\pi_{l}$ raises alerts for patients with risk scores greater than $T^{*}$. However, we do not know what patient outcomes are when $\pi_{l}$ raises alerts for patients with risk scores between $T^{l}$ and $T^{*}$.
@@
$Y(A=a,M=m_{i})\leq Y(A=a,M=m_{j})$
Assumption 3.1 says that improvements in model performance do not harm patient outcomes, given a fixed action. For instance, in the context of Example 1, we might expect that, for a given patient, having an alarm raised by a high-performance model would not lead to worse outcomes than if that alert had been raised by a model with frequent false alarms. Note that this assumption is stated with a fixed action $A=a$ and does not imply that improving performance alone is guaranteed to improve outcomes – a change in model performance is generally associated with a change in outputs, which may have its own effect on outcomes. In Section 5, we give a case where improved overall performance (accuracy) is associated with worse outcomes. Assumption 3.1, however, may not always hold; for instance, clinicians may begin paying less attention to patients that receive a low risk score from the DST, even if it is wrong, as their trust in the system increases. Such a scenario would violate Assumption 3.1. To address this, we propose a method for falsifying Assumption 3.1 below.
@@
implies that Assumption 3.1 is false. The proof for Assumption 3.1, along with all other proofs, is given in Appendix D. While it is not possible to guarantee that Assumption 3.1 is true in general (over all possible models), it has observable implications in an RCT that we can check. In particular, this result suggests a simple hypothesis test that we can use to falsify Assumption 3.1: compare two empirical means in the data and check if outcomes are lower under $\pi_{2}$ than under $\pi_{1}$ on those cases where $\pi_{1}$ and $\pi_{2}$ agree on their actions.
In our discussions thus far, we have considered the control arm to be just another policy. However, this framework creates practical difficulties when considering the model performance of a control arm. For instance, suppose the relevant metric for model performance is the false positivity rate (as in Example 1); then, $M$ is not clearly defined because the control arm never raises alerts. Alternatively, if the relevant metric were model accuracy under no deferral (as in Example 2), then the performance of the control arm would be similarly undefined. One way of resolving this tension is to presume the existence of a “neutral” action (e.g., not raising an alert, or deferring to clinicians).
@@
Assumption 3.2 is a sufficient condition for leveraging data from the control arm of an RCT in our setting, and it implies that, when a model output is “neutral” (e.g., no alert in Example 1, or deferral in Example 2), decision-makers act as they would if no model were deployed. However, note that there may not always exist a “neutral” output, especially if decision-makers tend to pay attention to model performance for all possible model outputs [^4] While all of our results make use of Assumption 3.2, they can also be re-written to hold if Assumption 3.2 is false by re-defining $a_{0}$ as some placeholder model output that is never observed under any model (including $\pi_{e}$), such that indicators like $\mathbf{1}\{\pi_{e}(x)=a_{0}\}$ are always zero.. Again, we propose a method for falsifying Assumption 3.2 below.
@@
Consider an evaluation policy $\pi_{u}(x)=\mathbf{1}\{r(x)>T^{u}\}$ where the threshold for alerting $T^{u}>T^{*}$ is greater than the threshold used in the original trial and where the performance of $\pi_{u}(x)$ (e.g., the precision) is greater than that of the original trialed policy $\pi_{1}$. Fig. 3 gives a visual illustration of such a policy. In this scenario, under Assumptions 3.1 and 3.2, we can intuitively infer a lower bound on the policy value of $\pi_{u}$ using the outcomes of both (a) patients with $r(x)\leq T^{u}$ who did not receive alerts in the trial (either because they were in the control arm or because $\pi_{1}$ did not raise alerts), and (b) those patients with $r(x)>T^{u}$ who did receive alerts under $\pi_{1}$.
While these assumptions are sufficient in some scenarios, they do not yield meaningful bounds when a new policy takes actions (i.e., a new model produces outputs) on a given case that was never seen for similar cases in the RCT.
@@
Consider the evaluation policy $\pi_{l}(x)=\mathbf{1}\{r(x)>T^{l}\}$ where the threshold for alerting $T^{l}<T^{*}$ is less than the threshold used in the original trial. Regardless of the performance of $\pi_{l}$ in this scenario, even under Assumptions 3.1 and 3.2, we have no way to infer outcomes under $\pi_{l}$ for the individuals where $r(x)\in[T^{l},T^{*}]$. These correspond to a set of “never alerted” individuals where $\pi_{l}$ raises an alert but where neither the control arm nor the trialed policy $\pi_{1}$ raised an alert.
@@
Aside: Why require the performance assumption? We pause to reflect on the importance of the assumption (implicit in Assumption 2.1) that implies that our choice of model $\Pi$ impacts outcomes, not only through the outputs $A$, but also through model performance $M$. Broadly speaking, this assumption is not only intuitive from a real-world perspective, but it also has the welcome side-effect of ruling out nonsensical conclusions about trial design. For instance, there are trivial ways to satisfy the condition that, for every value of $\mathcal{X}$, there exists some model in the trial that matches the output of $\pi_{e}$. In the context of Example 1, for instance, one could trial an alerting system that simply always raises alerts for every patient, alongside a control arm that never raises alerts. Then, the requirement that we observe what happens to patients both under no alerts and under alerts for each $x\in\mathcal{X}$ would be satisfied, eliminating any challenges related to coverage. The assumption that model accuracy $M$ impacts outcomes gives a formal rationale for why this type of trial design is nonsensical: The observed impact of this “always alert” policy would likely be minimal, or even harmful, compared to never raising alerts due to the negative impact of extremely poor accuracy.
We will shortly present our main result: Under our data-generating process (Assumption 2.1) and the assumptions above (Assumptions 3.1, 3.2 and 3.3), we can compute tight bounds on expected outcomes under any proposed model $\pi_{e}$. First, however, we will define some useful notation for conveying our results, which builds upon the intuition above.
@@
[Policy/Model Sets]definitionPartitionsDef For each value of $x\in\mathcal{X}$, we define the sets of trialed policies/models (possibly none) that agree with $\pi_{e}(x)$ and subsets of this set based on the performance characteristics of those trialed models [^5] All these sets are defined with respect to the model $\pi_{e}$ and could be written more precisely with $\pi_{e}$ as an argument instead of in the superscript (e.g., $\mathbf{\Pi}(x,\pi_{e})$) but we use the superscript notation for conciseness..
$\mathbf{\Pi}^{e}(x)$$\coloneqq\{\pi\in\varPi\mid\pi(x)=\pi_{e}(x)\}$
$\mathbf{\Pi}^{e}_{\leq}(x)$$\coloneqq\{\pi\in\varPi\mid\pi(x)=\pi_{e}(x),f_{M}(\pi)\leq f_{M}(\pi_{e})\}$
$\mathbf{\Pi}^{e}_{\geq}(x)$$\coloneqq\{\pi\in\varPi\mid\pi(x)=\pi_{e}(x),f_{M}(\pi)\geq f_{M}(\pi_{e})\}$
We also further define subsets of $\mathbf{\Pi}^{e}_{\leq}$ and $\mathbf{\Pi}^{e}_{\geq}$ that contain only the next-worst or next-best performing model [^6] Where relevant, we use the convention that $\operatorname*{arg\,min}_{\pi\in\varnothing}(f_{M}(\pi))=\varnothing$..
$\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)$ $\coloneqq\operatorname*{arg\,max}_{\pi\in\mathbf{\Pi}^{e}_{\leq}(x)}f_{M}(\pi),$
$\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)$ $\coloneqq\operatorname*{arg\,min}_{\pi\in\mathbf{\Pi}^{e}_{\geq}(x)}f_{M}(\pi)$
Remark 3.1.
To summarize our notation related to deployed and new models, we use $\varPi$ to denote the space of all deployed models in the RCT, $\Pi$ to refer to the variable representing the deployed model, $\pi$ to denote a specific deployed model, $\pi_{e}$ to denote the new model we are evaluating, and $\Pi$ in boldface with the superscript $e$, such as $\mathbf{\Pi}^{e}(x),\mathbf{\Pi}^{e}_{\leq}(x),\tilde{\mathbf{\Pi}}^{e}_{\leq}(x),\mathbf{\Pi}^{e}_{\geq}(x),$ and $\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)$, to denote functions that take a value of $x\in\mathcal{X}$ as input and return a set of models satisfying various criteria.
Using Assumption 3.3, we can give precise lower and upper bounds on the performance of a new model $\pi_{e}$.
{restatable}
theoremBoundsWithControl Given the data generating process in Assumption 2.1, and under Assumptions 3.3, 3.1 and 3.2, the policy value of a model / policy $\pi_{e}$ is bounded as